OpenStudy (anonymous):

sinx/1+cosx =1-cosx/sinx identities

myininaya (myininaya):

i think she means \[\frac{sinx}{1+cosx}=\frac{1-cosx}{sinx}\] is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cross multiply. add cos^2x on both sides

OpenStudy (anonymous):

Cross multiply :)

myininaya (myininaya):

i dont think you can do when you are trying to show something is something

myininaya (myininaya):

\[\frac{sinx}{1+cosx}*\frac{1-cosx}{1-cosx}=\frac{sinx*(1-cosx)}{1-\cos^2x}=\frac{sinx*(1-cosx)}{\sin^2x}\]

myininaya (myininaya):

=\[\frac{1-cosx}{sinx}\]

OpenStudy (anonymous):

and what would be the domain restriction on that?

myininaya (myininaya):

you don't want the bottom to be zero find when sinx=0 and whenever that happens thats what you want to exclude from your domain